OpenStudy (fantalux):

Which equation can be used to find the two numbers whose ratio is 4 to 3 and that have a sum of 42?

OpenStudy (anonymous):

x/y = 4/3 x=4/3y x+y = 42

OpenStudy (fantalux):

ty!